# RadMamba: Efficient Human Activity Recognition through Radar-based Micro-Doppler-Oriented Mamba State-Space Model

### v1

@@ III-C Issue 3: Challenges for Projection Layers in Capturing Doppler Patterns
Within the Mamba block, as shown in Fig. 1 top, the input $\mathbf{X_{p}}$ first undergoes layer normalization to yield $\mathbf{X}_{proj}$, which is then split into two branches via linear projections P1 and P2:
$\mathbf{Z}$ $=\mathbf{W}_{\textit{P1}}\mathbf{X}_{proj}+\mathbf{b}_{\textit{P1}},$ (2)
$\mathbf{X}_{fw}$ $=\mathbf{W}_{\textit{P2}}\mathbf{X}_{proj}+\mathbf{b}_{\textit{P2}},$ (3)
where the $\mathbf{W}$ and $\mathbf{b}$ terms are the weight matrices and bias vectors, respectively.
The projection operations are expected to extract Doppler frequency patterns over time and enhance the average cross-correlation between patches. This enhancement occurs because a high-performing classifier should recognize all Doppler-aligned vectors of a specific activity as a cohesive and highly relevant unit. However, the averaged cross-correlation between patches shows limited improvement from conventional projections P1 and P2. For instance, in Dataset CI4R [31], the average patch cross-correlation is 0.008 before projections 1 and 2. After projections 1 and 2, they rise to 77.9 and 0.4, respectively. In contrast to the improvements achieved through our proposed Conv1D projection (detailed later in Section IV-C), the linear layers in the conventional ViM architecture demonstrate limited capability to enhance cross-correlation between patches and consistently identify all time bins as components of the same action.
The tensors $\mathbf{X}{fw}$ and $\mathbf{X}{bw}$ resulting from the projections P1 and P2 are then passed through SSM branches. The outputs from both directions are combined using a gating mechanism modulated by the SiLU-activation to produce the gated output $\mathbf{Y}{sum}$. This gated output passes through projection layer P3, which is linear in the conventional ViM, and is added to the input via a residual connection:
$\mathbf{Y}_{out}=\mathbf{W}_{\textit{P3}}\mathbf{Y}_{sum}+\mathbf{b}_{\textit{P3}}+\mathbf{X}_{proj},$ (4)
Before and after projection 3, again using Dataset CI4R as an example, the average patch cross-correlation is 160.1 and 75.5, respectively. Essentially, the averaged cross-correlation between patches is even decreased by the conventional projection layer 3, posing a challenge to achieve good classification performance.
In the analysis of Issue 3, to compute the averaged patch cross-correlation, we first trained a model to its optimal performance, then duplicated and froze its parameters. By re-feeding the data into the frozen model, we calculated the averaged patch cross-correlation during inference. This averaged patch cross-correlation also accounts for different batches (classes). For example, with $\mathbf{X}{proj}\in\mathbb{R}^{B\times N\times D}$, each patch is labeled with its index $n$ in $\mathbf{X}{seg}$. The averaged patch cross-correlation is defined as:
@@
IV Proposed RadMamba
To overcome the aforementioned challenges and issues, we propose a micro-Doppler-oriented RadMamba architecture, illustrated in Fig. 2.
Figure 2: The schematic of the proposed RadMamba architecture and its novel solutions for radar micro-Doppler-based classification.
These modifications to the conventional approach include channel fusion and down-sampling, Doppler-aligned segmentation, and modified Mamba block rMamba.
IV-A Solution 1: Channel Fusion and Downsample
To tackle the issue of many similar and empty patches caused by high sparsity inherent in micro-Doppler signatures input $\mathbf{X}\in\mathbb{R}^{C\times H\times W}$, we introduce a preprocessing pipeline prior to the segmentation stage, as depicted in Fig. 2 Chan-DS: Channel Fusion and Downsample.
$\mathbf{X_{c-d}=\textit{Chan-DS}(\mathbf{X})}$ (7)
where $\mathbf{X}\in\mathbb{R}^{C_{c-d}\times H_{c-d}\times W_{c-d}},H_{c-d}<H,W_{c-d}<W$ is the input micro-Doppler signature after the channel fusion and downsample for each batch. The architecture of the Chan-DS block is as follows:
@@ IV-A Solution 1: Channel Fusion and Downsample
where $L=1,2$. The Two-Dimensional (2D) convolutional layer and batch normalization layer mix channel information and extract preliminary features from the input tensor $\mathbf{X}$. A 2D max pooling layer and a One-Dimensional (1D) max pooling layer reduces spatial dimensions. Optionally, an additional 1D average pooling layer with a kernel height as one and kernel width larger than 1, for example (1, 4), further downsamples the time dimension if the data sparsity along the time axis, influenced by the Pulse Repetition Interval (PRI) of the radar and pre-processing to generate the input, is excessive.
In combination with the Doppler-Aligned Segment explained in the next subsection IV-B, this method avoids the issue of having many similar and empty patches of data. Moreover, the down-sampling decreases the available size for patch segmentation, which enhances the energy-efficient design.
IV-B Solution 2: Doppler-Aligned Segment and Patch Embedding
To maintain the temporal and Doppler location information embedded in micro-Doppler signatures, we propose a new patch segmentation strategy, as shown in Fig. 2 Solution 2 Doppler-Aligned Segment, DA-Seg:
$\mathbf{X_{seg}}=\textit{DA-Seg}(\mathbf{X}_{c-d})$ (8)
This approach ensures that the temporal evolution of state sequence $N$ remains as $[t_{0},t_{1},\cdots,t_{J}]$ and continuous as time sequence $W$. When the data is sent to the SSM, as depicted in Fig.2 Solution 2 left, this method avoids the loss of critical temporal information that conventional square patches of the ViM might face.
Moreover, this segmentation approach results in a pixel sequence dimension $CH_{seg}W_{seg}=H_{c-d}$ of $\mathbf{X}^{\prime}_{seg}$ that preserves the frequency pattern as $[f_{0},f_{1},\cdots,f_{h\_seg}]$, during Flatten operations within each Doppler bin, as shown in Fig.2 Solution 2 left, as:
$\mathbf{X^{\prime}_{seg}}=\textit{Flatten}(\mathbf{X_{seg}})=\mathbf{X_{seg}}$ (9)
IV-C Solution 3: Convolution Projections in rMamba Blocks
Finally, to increase the capability of projection layers to capture the Doppler patterns, we replace the conventional linear projection layers within the conventional Mamba block with 1D convolutional layers, as highlighted in Fig. 2 rMamba Block, in the blue boxes on the top-right part of the figure.
Within the rMamba block, the linear projections P1 and P2 are replaced by Conv1D layer as:

### v3

@@ III-C Solution 3: Convolution Projections in CP-Mamba Blocks
Within the conventional Mamba block (Fig. 1), the input $\mathbf{X}_{p}$ is first normalized to obtain $\mathbf{X}_{proj}$, which is then split into three paths:
$\mathbf{Z}$ $=\text{Projection}_{P1}(\mathbf{X}_{proj}),$ (9)
$\mathbf{X}_{fw}$ $=\text{Projection}_{P2}(\mathbf{X}_{proj}),$ (10)
$\mathbf{X}_{bw}$ $=\text{Flip}(\mathbf{X}_{fw},dim=\text{temporal dim.}),$ (11)
where $\mathbf{Z},\mathbf{X}_{fw},\mathbf{X}_{bw}\in\mathbb{R}^{N\times dim}$, and Flip() reverses the temporal dimension of $\mathbf{X}_{fw}$.
In the conventional Mamba block, $\text{Projection}_{P1}$ and $\text{Projection}_{P2}$ are linear layers. In contrast, the CP-Mamba block replaces them with 1D convolutions using kernel size $(dim,dim,3)$. Next, $\mathbf{X}_{fw}$ and $\mathbf{X}_{bw}$ are processed by the two SSM branches:
$[\text{Conv1D}{fw/bw},\text{LayerNorm}{fw/bw},\text{SSM}{fw/bw}],$ (12)
where $\text{Conv1D}{fw/bw}$ has weight shape $(dim,dim,1)$ and operates along the sequence dimension $(N,dim)$.
The two directional outputs are gated by the Sigmoid Linear Unit (SiLU)-activated $\mathbf{Z}$ and summed:
$\mathbf{Y}^{\prime}{fw/bw}$ $=\mathbf{Y}{fw/bw}\odot\text{SiLU}(\mathbf{Z}),$ (13)
$\mathbf{Y}{sum}$ $=\mathbf{Y}^{\prime}{fw}+\mathbf{Y}^{\prime}{bw},$ (14)
where $\odot$ denotes element-wise multiplication. Finally, a projection P3 and residual connection produce the output:
$\mathbf{Y}_{out}=\text{Projection}_{P3}(\mathbf{Y}_{sum})+\mathbf{X}_{proj},$ (15)
In the conventional Mamba block, the projection layer P3 is linear. In our CP-Mamba, P3 is replaced with a Conv1D, where $\text{Conv1D}_{P3}$ has weights $(dim,dim,1)$ and restores the sequence shape to $(N,dim)$.
TABLE I: Averaged cross-correlation between patches before and after each projection layer.
Linear Projection Layers
in Conventional Mamba
Convolutional Projection Layers
in Our Proposed CP-Mamba
Projection At Input At Output At Input At Output
1 0.008 77.9 0.002 1065.7
2 0.008 0.4 0.002 16.1
3 160.1 75.5 245.3 356.4
Ideally, projection layers should improve cross-correlation between patches, since a strong classifier should treat Doppler-aligned patches of the same activity as a coherent unit. Linear projections, however, increase correlation only marginally (P1, P2) or even reduce it (P3), limiting classification performance. The convolutional projections in CP-Mamba consistently strengthen cross-correlation, making patches easier to identify as belonging to the same action and improving Doppler-pattern sensitivity.
Tab. I reports the averaged patch cross-correlation using the Dataset CI4R [26]. For conventional linear projections, the averaged correlation before projections 1 and 2 is 0.008, rising to 0.77 and 0.4, respectively. Before and after projection 3, it drops from 160.1 to 75.5. In contrast, our convolutional projections greatly enlarge the correlation gap (e.g., from 77.9 to 1065.7 for P1), indicating superior ability to aggregate time bins of the same activity.
To compute the averaged patch cross-correlation, we train the model, freeze the parameters, and refeed the data. For $\mathbf{X}_{proj}\in\mathbb{R}^{B\times N\times D}$, each patch is indexed by $n$. The averaged cross-correlation is defined as:
@@
III-D Selective State-Space Modeling for Micro-Doppler Sequences
Figure 3: The architecture of the State Space Model in RadMamba.
The SSM is a significant layer in CP-Mamba that enables efficient sequence modeling for the dynamic nature of micro-Doppler signatures. As shown in Fig. 3 (a), the discrete SSM is defined as:
$\mathbf{y}[n]$ $=\mathbf{C}\mathbf{h}[n]+\mathbf{D}\mathbf{x}[n],$ (19)
$\bar{\mathbf{A}}$ $=\exp(\boldsymbol{\Delta}\cdot\mathbf{A}),$ (20)
$\bar{\mathbf{B}}$ $=\left(\int_{0}^{\boldsymbol{\Delta}}\exp(\mathbf{A}\tau)\,d\tau\right)\mathbf{B},$ (21)
The input of the SSM branches in the proposed CP-Mamba block are tensors $\mathbf{X}_{fw/bw}\in\mathbb{R}^{N\times dim}$; after Conv1D and LayerNorm in SSM branch, this is termed as $\mathbf{X^{\prime}}_{fw/bw}\in\mathbb{R}^{N\times dim}$.
Therefore, the discretized state-transition tensor $\mathbf{\bar{A}}$} has dimension $N\times dim\times dim_{s}$, and the input tensor $\mathbf{\bar{B}}$} has dimension $N\times N\times dim_{s}$. Here, $N$ equals the time dimension in the micro-Doppler signatures. The input tensor $\mathbf{\bar{B}}$ captures how the Doppler-aligned input vector influences the state $\mathbf{h}[n]$, whereas the state-transition tensor $\mathbf{\bar{A}}$} captures how the previous state $\mathbf{h}[n-1]$ contributes to $\mathbf{h}[n]$. Moreover, the matrices $\mathbf{C}\in\mathbb{R}^{N\times dim_{s}}$ translate the current state $\mathbf{h}[n]$ to the output. The skip matrices $\mathbf{D}\in\mathbb{R}^{dim}$ add weighted input directly to the output of SSM, which helps the recurrent part focus on the temporal evolution.
$\mathbf{B}$ $=\mathbf{W}_{\mathbf{B}}\mathbf{X},$ (22)
$\mathbf{C}$ $=\mathbf{W}_{\mathbf{C}}\mathbf{X},$ (23)
where $\mathbf{W}_{\mathbf{B}}$ and $\mathbf{W}_{\mathbf{C}}$ are weights of dimension $(dim,dim_{s})$. The $\boldsymbol{\Delta}$ computation involves a linear layer with weights $(dim,\text{dt\_rank})$ followed by a projection to $(N,dim)$ via weights $(\text{dt\_rank},dim)$, where dt_rank is a hyperparameter controlling the rank of the dynamic tensor. The Softplus activation ensures positive step sizes, implemented as $\text{Softplus}(x)=\log(1+\exp(x))$.
III-E Overall Architecture
To summarize the method, the proposed RadMamba pipeline in pseudo-code is presented in Algorithm I. The patch embedding in PEmb follows the method in  [18], and the position embedding method is a sinusoidal position encoding. [34].
Algorithm 1 Proposed RadMamba Pipeline
Parameters: dimdim, $dim_{s}$, dt_rankdt\_rank
1:Channel Fusion and Downsample:
[…]
3:Doppler-Aligned Segmentation:
[…]
5:Patch and Position Embedding:
[…]
7:CP-Mamba Block:
[…]
9:Classify:
[…]
IV Open-source Datasets and Preprocessing for Radar-based HAR
To evaluate the efficacy of our proposed RadMamba framework, we utilize three distinct radar datasets. This section describes these datasets, namely the CW non-continuous Dataset DIAT, FMCW non-continuous Dataset CI4R, and FMCW continuous Dataset UoG20, along with their respective pre-processing steps. More details on these datasets are summarized in Tab. II. Here, the word ‘continuous’ refers to samples of data where sequences of human activities are continuously performed one after the other. Examples of micro-doppler signature after pre-processing are shown in Fig. 4.
TABLE II: Overview of radar-based human activity recognition datasets used in this work.
Datasets DIAT [25] CI4R [26] UoG20 [27]
Continuous No No Yes
Classes 6 11 6
Sensor X-band CW IWR1443 FMCW FMCW
Freq.(GHz) 10 77 5.8
[…]
BW(MHz) N/A 750 400
Length per sample (s) 3 20 35
Train Samples 2646 528 39
Test Samples 759 132 6
Sliding Window N/A N/A 224
Figure 4: Micro-Doppler Signature Example of (a) Dataset DIAT (b) Dataset CI4R (c) Dataset UoG20
IV-A Dataset DIAT: X-Band CW Radar with Discrete Activity Instances [25]
This dataset was collected using an X-band CW radar operating at a frequency of 10 GHz. It encompasses data from 30 human subjects performing ‘suspicious’ activities at ranges spanning 10 meters to $\pm$0.5 kilometers over 3-second durations, resulting in an imbalanced set of 3,780 spectrogram samples of shape $(3,224,224)$ across six activity classes. Preprocessing adheres to the methodology outlined in [25]. To facilitate a fair comparison with previous works, the dataset is partitioned into 70% training, 10% validation, and 20% test sets within each class, maintaining class-specific balance for reliable performance evaluation.
IV-B Dataset CI4R: Short-Range FMCW Radar for Multi-Class Activity Recognition [26]
For the second dataset, experiments employ a Texas Instruments IWR1443 FMCW radar, operating at 77 GHz with a bandwidth of 750 MHz. It includes data from six participants, varying in age, height, and weight, performing 11 distinct activities and ambulatory gaits. Each participant executed 10 repetitions per activity, yielding 60 samples per class per sensor, as detailed in [26]. Preprocessing involves applying a 256-point short-time Fourier transform ( Short-Time Fourier Transform (STFT)) with a context window length of 256 to convert the raw radar signals into a 2D micro-Doppler representation of shape $(1,224,224)$. The dataset is split with an 8:2 training-to-test ratio, ensuring sufficient data for model optimization and evaluation.
IV-C Dataset UoG2020: Continuous FMCW Radar for Realistic HAR Sequences [27]
Representing a leap in complexity, this dataset was acquired using an FMCW radar operating at 5.8 GHz with a 400 MHz bandwidth. It comprises data from 15 participants (14 males, 1 female, aged 21–35) performing six basic activities within continuous 35-second sequences. These activities are organized into three distinct sequential orders, with participants granted autonomy in selecting transition points between consecutive activities and determining trajectories for translational tasks, as described in [27]. The resulting micro-Doppler signatures are preprocessed into $(1,240,224)$ 2D frames, with data from two participants designated as the test set and data from the remaining 13 participants used for training, providing a challenging benchmark for continuous sequence modeling.
V Experimental Setup
V-A Network Structure
TABLE III: Network Architecture of Models Used in This Work
Model Swept Hidden Size (dimdim for RadMamba)
ResNet6 (Customized) 5, 7, 12, 20, 30
Bi-LSTM 1, 2, 4, 10, 20, 28, 34
CNN-LSTM 2, 4, 6, 16, 28, 34, 46
CNN-Bi-GRU 1, 2, 4, 6, 16, 28
RadMamba Dataset DIAT 8, 16, 32, 64, 80
RadMamba Dataset CI4R 8, 16, 32, 64, 80, 96, 128, 160
RadMamba Dataset UoG20 8, 16, 20, 24, 32
The classification performance of RadMamba is evaluated against a suite of established models, selected for their proven efficacy across CNN, RNN, and hybrid architectures: customized ResNet6 [35], Bi-LSTM [27], CNN-LSTM [16], and CNN-Bi-GRU [17]. To explore the impact of model scaling, we conducted a model size sweep experiment, varying the hidden dimensions as shown in Tab. III. These sizes are chosen to capture the full spectrum of performance trends as model capacity increases, providing insights into scalability and efficiency.
The customized ResNet6 is derived from ResNet18 by preserving the first five layers and the final classifier while maintaining its channel ratio between consecutive channels. The minimum channel width is tunable to target low-FLOP regimes, enabling fair comparisons under limited computation conditions. Although the CNN-Bi-GRU retains a larger parameter budget, its fusion strategy may not generalize consistently across datasets.
All four baselines (ResNet6, Bi-LSTM, CNN-LSTM, and CNN-Bi-GRU) are retrained from scratch in a unified PyTorch pipeline, using consistent dataset splits, preprocessing, optimization settings, and batch=1 memory profiling. For transformer and SSM competitors, FML-ViT [19], LW-ViT [33], and ActivityMamba [22], we report their best published small-scale configurations. Following their protocols, we adopt the same training settings for FML-ViT and ActivityMamba, but use a different input size than LW-ViT: their preprocessing produces 656×656 inputs on the CI4R dataset, whereas our pipeline uses 224×224.
TABLE IV: Specified Network Architecture Variations of RadMamba
Datasets $dim_{s}$ dt_rankdt\_rank Chan. Fusion DS Redu. Factor
Dataset DIAT 1 2 2, (16, 3, 3) (2, 2)
Dataset CI4R 4 0 1, (1, 3, 3) (2, 8)
Dataset UoG20 16 4 1, (1, 3, 3) (2, 32)
For RadMamba, the configurations of channel fusion, downsampling, and SSM parameters are dataset-specific to optimize micro-Doppler processing (Tab. IV). The ‘channel fusion’ column specifies [L, ($C_{c-d},H_{k},W_{k}$)], where L is the number of channel fusion blocks and the $C_{c-d},H_{k},W_{k}$ are channel, kernel height, kernel width of Conv2D layer, respectively. The ‘DS Redu. Factor’ denotes the reduction factors ($H/H_{c-d}$, $W/W_{c-d}$).
Hyperparameter selection follows two principles:
  • Bayesian optimization: Because RadMamba is highly configurable, we perform a Bayesian-optimization-based search over a constrained space. Since dimdim predominantly determines the sizes of most weight matrices, the number of channels in Chan-DS, $dim_{s}$, and $dt_{rank}$ were selected to minimize model complexity across different dimdim.
  • Physical perspective from micro-Doppler signature (Temporal downsampling and Reduction Factor): The first downsampling factor of $2\times 2$ in Chan-DS matches the fact that most Doppler energy appears in the central part of the spectrum. With horizontal reduction factors of $8$ and $32$ for the CI4R and UoG20 datasets (as shown in Tab. IV), each Doppler vector spans approximately 640  ms. Empirically, this temporal interval provides an optimal balance for both FMCW datasets.
V-B Training Hyperparameter
Training is conducted using the Adaptive Moment Estimation with decoupled Weight decay (AdamW) optimizer [36, 37], paired with the ReduceLROnPlateau scheduler, which adjusts the learning rate adaptively based on test performance. For Dataset DIAT, the initial learning rate is set to $1\times 10^{-4}$, with a batch size of 16, and training proceeds for 50 epochs. For Dataset CI4R, RadMamba employs an initial learning rate of $5\times 10^{-3}$, while the baseline models use $1\times 10^{-4}$, with a batch size of 16 over 100 epochs, reflecting the dataset’s increased complexity and sample size. For Dataset UoG20, the initial learning rate is $5\times 10^{-5}$, with a larger batch size of 256 and 50 epochs, processing continuous sequences with a frame length of 224 and a stride of 1 to effectively capture temporal dependencies. All experiments are executed on a single NVIDIA RTX 4090 GPU.
To account for training variability, all models, including RadMamba, are trained and evaluated ten times with distinct seeds. The following analysis reports the mean and standard deviation of performance metrics from the model size swept experiment.
V-C Evaluation Settings
Inference time and memory profiling are conducted on an NVIDIA Jetson Orin Nano Developer Kit with PyTorch 2.8.0 and CUDA 12.6. The kit is equipped with a 6-core Arm Cortex-A78AE CPU and an NVIDIA Ampere Architecture GPU with unified off-chip 8 GB 128-bit LPDDR5 memory. The evaluation follows a standardized protocol: each model undergoes 10 warm-up iterations followed by 100 measurement iterations, with inference performed at a batch size of 1. Peak memory (Peak Mem.) is measured using torch.cuda.memory_allocated(device), which returns the maximum allocated memory across all 100 inference iterations. CUDA synchronization with the ARM CPU is enforced before and after each forward pass to ensure accurate timing measurements.
VI Results and Discussion
VI-A Comparison with Previous Works
TABLE V: Mean and Standard Deviation of Classification Accuracy Performance across Seed 0 to 9 of Different NN-based RadHAR Models Evaluated With Three Different Radar Settings alongside Their Model Size and Floating-point Operations (#\#FLOP.)
Dataset DIAT [25]
(CW, Non-continuous)
Dataset CI4R [26]
(FMCW, Non-continuous)
Dataset UoG20 [27]
(FMCW, Continuous)
Classifiers
#params
#\#FLOP a
Accuracy
Peak Mem
Inf. Time
#params
#\#FLOP. a
Accuracy
Peak Mem.
Inf. Time
#params
#\#FLOP a
Accuracy
Peak Mem.
Inf. Time
@@ VI-A Comparison with Previous Works
FML-ViTc [19] - - - - - $\sim$2700 169.0 $\sim$92.0 - - - - - - -
LW-VITc [33] - - - - - 769 $\sim$2410 $\sim$92.0 - - - - - - -
ActivityMambac [22] $\sim$8700.0 $\sim$1220.0 99.8 - - - - - - - - - - - -